3(4+x)=18
2x-1=3/2
1
Expert's answer
2012-07-24T07:25:42-0400
1.
3(4+x)=18 <=> 12+4x=18 <=> 4x=6 <=> x=3/2

2.
2x-1=3/2
2x=1+3/2=5/2
x=5/4